if 2 is subtracted from a number and this difference is tripled, the result is 6 more than the number. find the number? im really having a hard time with this one. so far this is what i have. 2x-3=6+x. please let me know if i fomatted the problem correctly? thanks.
Answer by askmemath(368)

Let the number = X
2 is subtracted => X-2
Tripled = 3X-6
6 more than the number
3X-6 = X+6
Adding 6 on both sides
3X = X+12
Subtracting X on both sides
2X = 12
Dividing by 2 throughout
X = 6
Check:
2 subtracted = 6-2 = 4
Tripled = 12
which is 6 more than the original number
which means our answer is correct Yay!!!